OpenStudy (ashleytull):

What is the domain of the function f(x) = x2 + 3x + 5?

OpenStudy (brainlesslonelygirl):

3/2?

OpenStudy (mathmale):

Be certain you understand what "domain" means in this context. Rule: The domain of all polynomials is the set of all real numbers. Is f(x) = x^2 + 3x + 5 a polynomial, or not? f(x) = x2 + 3x + 5 should be written as f(x) = x^2 + 3x + 5. x^2 represents "the square of x." x2 is vague.

likeabossssssss (likeabossssssss):

\(\bbox[5pt,midnightblue]{\LARGE\color{yellow}{\star ^\bigstar Welcome~to~~os~ ^\bigstar \star}}\)